If Chameli had ₹600 left after spending 75% of her money, how much did she have in the beginning? 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Let the amount of money which Chameli had in the beginning be x.
Given that, after spending 75% of₹x, she was left with₹600.
∴ (100 − 75)% of x=₹600
Or, 25 % of x =₹600
25/100 × 𝑥=₹600
x = 600×4 = 2400
∴ Chameli had₹2400 in the beginning

Thanks